OpenStudy (anonymous):

A is the midpoint of XY Find: XA Find: AY and XY X--^3x---A---^5x-6--Y

OpenStudy (antonio_xx2):

XA and AY bot have equal measurement. And their sum is equal to XY.

OpenStudy (anonymous):

XA = 1/2 of XY AY = 1/2 of XY XY = XY